The problem says for me to solve but I don't even know how to begin. I know I need an LCM but what is it, 3x? Then what, do I divide both sides by the 3x? WHen I did that I came up with 6 = 21 + x
solve for x: 6-21=x
-15=x
Is this right?
2/x = 7/x - 1/3

Answer by stanbon(57354)

2/x = 7/x - 1/3
----
Multiply both sides by 3x to get rid of the denominators.
6 = 21 - x
x = 21-6
x = 15
=================
Cheers,
Stan H.